A species of horned toads has a (starting) population size of 5000 individuals. Every month, 400 individuals are born. However, you cannot assess the number of individuals that die every month because after death the animals are susceptible to a fungus that quickly degrades their carcasses. You can measure, though, that the population is declining over time, with the population size decreasing by 250 individuals every month (i.e., deltaN/deltaT). Calculate the total number of deaths per month and the per capita death rate. (Use 5,000 as your value for N.) Show your work.
Answer #1

Number of horned taods in given population (N) = 5000

Population decline over time (dN/dT) = 250 / month

Intrinsic rate of increase in population (rN) = Birth - death

                                                                      = 400 - 250 = 150

r = 250 / 150 = 1.66

Per capita death rate = D/N = number of deaths per month / total number of population = 250 / 5000 = 0.05